# Preference-Conditioned Gradient Variations for Multi-Objective Quality-Diversity

### v1

@@ Abstract
In a variety of domains, from robotics to finance, Quality-Diversity algorithms have been used to generate collections of both diverse and high-performing solutions. Multi-Objective Quality-Diversity algorithms have emerged as a promising approach for applying these methods to complex, multi-objective problems. However, existing methods are limited by their search capabilities. For example, Multi-Objective Map-Elites depends on random genetic variations which struggle in high-dimensional search spaces. Despite efforts to enhance search efficiency with gradient-based mutation operators, existing approaches consider updating solutions to improve on each objective separately rather than achieving desired trade-offs. In this work, we address this limitation by introducing Multi-Objective Map-Elites with Preference-Conditioned Policy-Gradient and Crowding Mechanisms: a new Multi-Objective Quality-Diversity algorithm that uses preference-conditioned policy-gradient mutations to efficiently discover promising regions of the objective space and crowding mechanisms to promote a uniform distribution of solutions on the Pareto front. We evaluate our approach on six robotics locomotion tasks and show that our method outperforms or matches all state-of-the-art Multi-Objective Quality-Diversity methods in all six, including two newly proposed tri-objective tasks. Importantly, our method also achieves a smoother set of trade-offs, as measured by newly-proposed sparsity-based metrics. This performance comes at a lower computational storage cost compared to previous methods.
Index Terms:
Quality-Diversity, Multi-Objective optimisation, MAP-Elites, Neuroevolution, Reinforcement Learning
I Introduction
@@
Figure 1: Left. Multi-Objective MAP-Elites repertoire. The feature space $\mathcal{C}\subset\mathbb{R}^{d}$ is tessellated into cells $\mathcal{C}_{i}$. A Pareto Front is stored in each cell. The aim of moqd algorithms is to fill each cell with solutions that are Pareto-optimal. Right. Overview of preference-conditioned policy gradient in the mome-p2c algorithm. By conditioning policy-gradients on updates solutions can be improved toward achieving different trade-offs of objectives (illustrated by blue arrows). By contrast, in mome-pgx, solutions are only updated to improve performance on each objective separately (illustrated by light blue arrows).
@@
Over recent years, Deep Reinforcement Learning (RL) has enabled breakthroughs in mastering games [1, 2] as well as continuous control domains for locomotion [3, 4] and manipulation [5]. These milestones have demonstrated the extraordinary potential of RL algorithms to solve specific problems. However, most approaches return only one highly-specialised solution to a single problem. In contrast, there is a growing shift in focus towards not just uncovering one single solution that achieves high rewards, but instead many solutions that exhibit different ways of doing so [6]. Within this context, Quality-Diversity (qd) algorithms [7] have emerged as one promising approach for tackling this challenge.
@@
In qd, the primary goal is to produce a variety of high-quality solutions, rather than to focus exclusively on finding the single best one. One motivation for qd algorithms is that, finding many solutions can provide availability of alternative, back-up solutions in the event that the highest-performing solution is no longer suitable. For example, in robotics, generating large collections of solutions has been shown to be helpful for addressing large simulation to reality gaps [8] and adapting to unforeseen damages [8, 9]. Alternatively, having multiple solutions can simply be used in order to promote innovation in the downstream task. In this context, qd has been used for creating diverse video game levels [10, 11] and generating building designs [12].
@@
Despite the growing traction of qd, most research in this field has focused on single-objective applications. However, multi-objective (mo) problems pervade many real-world domains, including engineering [13, 14], finance [15], and drug design [16] and many state-of-the-art mo algorithms originate from Evolutionary Algorithm community [17, 18, 19, 20].
@@
Recently, Multi-Objective MAP-Elites algorithm (mome) [21] marked the first attempt at bridging ideas from qd and mo optimisation. In moqd, the overarching goal is to identify a broad collection of solutions that exhibit diverse features and achieve distinct performances across multiple objectives. More specifically, given a feature space that is tessellated into cells, the aim is to find a collection of solutions within each cell which offer different trade-offs on each of the objectives (see Figure 1). As an example, consider the task of designing building sites. Within this context, it may be interesting to find different designs that vary in the number of buildings on the site. Then for each possible number of buildings, further options can be generated which present different trade-offs of ventilation and noise levels [12]. This approach equips end-users with a spectrum of viable options, thereby broadening their perspective on the array of feasible design possibilities.
@@
The mome algorithm demonstrated promising results in finding large collections of diverse solutions that balance multiple objectives. However, mome predominantly depends on random genetic variations that can cause slow convergence in large search spaces [22, 23, 24]. This renders it less suitable for evolving neural networks with a large number of parameters.
@@
Since the inception of the mome framework, several related works exploring the domain of moqd have emerged [25, 12, 26]. Among them, mome-pgx [25] builds upon the mome framework and was shown to achieve state-of-the-art performance on high-dimensional continuous control robotics tasks that can be framed as Markov Decision Processes. It uses crowding addition and selection mechanisms to encourage an even distribution of solutions on the Pareto front and employs policy-gradient mutations for each objective function in order to drive the exploration process toward promising regions of the solution space. However, the mome-pgx approach is not without its own set of challenges. Firstly, it employs separate actor-critic networks for each objective function, which can be resource-intensive and may not scale with an increasing number of objectives. Furthermore, although using policy gradient-based updates helps with exploration in high-dimensional search spaces, the approach in mome-pgx only considers improving solutions on each objective separately. However, in the context of multi-objective problems, the goal is often not just to maximise each objective independently but rather to find solutions which offer different trade-offs among them. In this way, if end users have different preferences regarding the relative importance of each objective, they have a range of solutions to choose from.
@@
In this paper, we address the limitations of mome-pgx by introducing a new moqd algorithm: Multi-Objective Map-Elites with Preference-Conditioned Policy-Gradient and Crowding Mechanisms (mome-p2c). Rather than using a separate actor-critic framework for each objective, mome-p2c uses a single preference-conditioned actor and a single preference-conditioned critic. Similar to mome-pgx, the actor-critic framework in mome-p2c can be used to provide policy-gradient mutations which offer efficient search space exploration for high-dimensional neural-network policies. However, as illustrated in Figure 1, by conditioning the actor and critic networks on a preference, policy-gradient updates can be used to improve solutions toward achieving a given weighting over the objectives, rather than improve solutions on each objective disjointly. Moreover, using a single preference-conditioned actor-critic framework rather than one per objective also reduces the memory costs and training costs associated with maintaining the separate actor-critic networks of mome-pgx.
@@
We show that mome-p2c outperforms or matches the performance of mome-pgx across six robotic control moqd tasks, including newly introduced tri-objective ones (see Section V-A). mome-p2c also outperforms mome-pgx on two newly introduced sparsity-based moqd metrics (see Section V-C) demonstrating that it is able to attain a smoother set of trade-offs than mome-pgx. The code for mome-p2c is fully containerised and available at Code hidden for anonymity, will be released upon acceptance..
II Background
II-A Quality-Diversity
@@
Quality-Diversity algorithms aim to discover collections of solutions that are both high-performing and diverse [27]. Similar to standard optimisation algorithms, a solution $\theta\in\Theta$ is assessed via a fitness function $f:\Theta\to\mathbb{R}$ that reflects its performance on the task. For example, consider the task of generating an image of a celebrity from a text prompt. In this case, the fitness of a solution could be the CLIP score [28] which measures the fidelity of an image to its caption that was used to generate it. However, an additional central component to qd algorithms, is the concept of the feature function $\Phi:\Theta\to\mathbb{R}^{d}$ that characterizes solutions in a meaningful way for the type of diversity desired [27]. The feature of a solution $\Phi(\theta_{i})$ is a vector that captures some of its notable characteristics, which is then consequently used to quantify its novelty relative to other solutions. In the image generation example, the feature could be the hair length or age of the subject in the photo [29]. In this example, the qd algorithm would then aim to generate images in which the subject has a diverse range of hair lengths and ages, and which closely obey the given text prompt used to generate it.
@@
One branch of algorithms in the qd family stems from the map-elites algorithm [30], which has gained prominence for its simplicity and effectiveness. map-elites operates by discretising the feature space into a grid-like structure, where each cell $C_{i}$ of the grid becomes a “niche” that can be occupied by a solution. Tessellating the feature space in this manner creates a systematic method for exploring of different niches within this space [31]. Each iteration of map-elites first involves selecting solutions from these niches, creating copies of them and mutating these copies to create new candidate solutions. Then, the fitness and features of the candidate solutions are evaluated, and they are added to the appropriate niches based on their fitness. If the cell corresponding to the new solution’s feature vector is unoccupied, the new solution is added to the cell. If the cell is occupied, but the evaluated solution has a higher fitness than the current occupant, it is added to the grid. Otherwise, the solution is discarded. This process continues for a fixed number of iterations, progressively populating the grid structure with an array of diverse, high-quality solutions.
@@
map-elites algorithms aim to maximise the total number of occupied cells at the end of the process and the performance of the solutions within each of them. Given a search space $\Theta$ and a feature space $\mathcal{C}$ that has been tessellated into $k$ cells $\mathcal{C}_{i}$, the map-elites objective, or QD-score [7] can be formally expressed as:
$\max_{\theta\in\Theta}\sum_{i=1}^{k}f(\theta_{i}),\,\,\text{where}\,\,\forall i,\Phi(\theta_{i})\in\mathcal{C}_{i}$ (1)
II-B Multi-Objective Optimisation
@@
Multi-Objective (mo) optimization provides an approach for addressing problems that involve the simultaneous consideration of multiple, often conflicting objectives $\textbf{F}=[f_{1},\dots,f_{m}]$. In mo problems, objectives often compete with each other, meaning that improving one objective typically comes at the expense of another. For example, in engineering, improving performance might increase cost, and vice versa. To navigate this landscape, the concept of Pareto-dominance is commonly employed to establish a preference ordering among solutions. A solution $\theta_{1}$ is said to dominate another solution $\theta_{2}$ if it is equal or superior in at least one objective and not worse in any other [32]. That is, $\theta_{1}\succ\theta_{2}$, if $\forall i:f_{i}(\theta_{1})\geq f_{i}(\theta_{2})\wedge\exists j:f_{j}(\theta_{1})>f_{j}(\theta_{2})$.
@@
Solutions that are not dominated by any other solutions are termed non-dominated. Given a set of candidate solutions $\mathcal{S}$, the non-dominated solutions of this set $\theta_{i}\in S$ collectively form a Pareto front, which represents the boundary of achievable trade-offs among objectives. The goal of mo optimisation is to find an approximation to the optimal Pareto front, which is the Pareto front over the entire search space $\Theta$.
Figure 2: Sets of solutions that form approximations to two Pareto fronts. The hypervolume of the outer solutions is larger as they achieve higher performance on the objectives. Likewise, the sparsity metric of the outer solutions will also be higher as they are more evenly spaced.
@@
There are two metrics, the hypervolume and sparsity metric (see Figure 2), that play pivotal roles in comprehensively assessing the quality and diversity of solutions within the Pareto front [33, 32]. The hypervolume $\Xi$ of a Pareto front $\mathcal{P}$, measures the volume of the objective space enclosed by a set of solutions relative to a fixed reference point $r$. This metric provides a quantitative measure of the quality and spread of solutions in the objective space and is calculated as [32, 33]:
$\Xi(\mathcal{P})=\lambda({\theta\in\Theta\,|\,\exists\,s\in\mathcal{P},s\succ x\succ r})$ (2)
@@
While the hypervolume metric quantifies the coverage of the objective space by solutions on the Pareto front, sparsity provides complementary information regarding the distribution and evenness of these solutions. It is calculated by evaluating the average nearest neighbour distance among solutions on the Pareto front, given by [32]:
$S(\mathcal{P})=\frac{1}{|\mathcal{P}|-1}\sum_{j=1}^{m}\sum_{i=1}^{|\mathcal{P}|-1}(\tilde{P_{j}}(i)-\tilde{P_{j}}(i+1))^{2}$ (3)
where $\tilde{P_{j}}(i)$ denotes the $i$-th solution of the list of solutions on the front $\mathcal{P}$, sorted according to the $j$-th objective and $|\mathcal{P}|$ denotes the number of solutions on the front. To ensure that the sparsity is not skewed due to different scales of each of the objectives, the objective functions must be normalised prior to calculating it.
@@
A low-sparsity metric indicates that solutions are well-dispersed through the objective space, highlighting the algorithm’s ability to provide diverse trade-off solutions. In contrast, a high-sparsity metric suggests that solutions are clustered in specific regions, potentially indicating that the algorithm struggles to explore and represent the full range of possible trade-offs.
II-C Multi-Objective Quality-Diversity Algorithms

### v2

@@ Abstract.
In a variety of domains, from robotics to finance, Quality-Diversity algorithms have been used to generate collections of both diverse and high-performing solutions. Multi-Objective Quality-Diversity algorithms have emerged as a promising approach for applying these methods to complex, multi-objective problems. However, existing methods are limited by their search capabilities. For example, Multi-Objective Map-Elites depends on random genetic variations which struggle in high-dimensional search spaces. Despite efforts to enhance search efficiency with gradient-based mutation operators, existing approaches consider updating solutions to improve on each objective separately rather than achieving desired trade-offs. In this work, we address this limitation by introducing Multi-Objective Map-Elites with Preference-Conditioned Policy-Gradient and Crowding Mechanisms: a new Multi-Objective Quality-Diversity algorithm that uses preference-conditioned policy-gradient mutations to efficiently discover promising regions of the objective space and crowding mechanisms to promote a uniform distribution of solutions on the non-dominated front. We evaluate our approach on six robotics locomotion tasks and show that our method outperforms or matches all state-of-the-art Multi-Objective Quality-Diversity methods in all six, including two newly proposed tri-objective tasks. Importantly, our method also achieves a smoother set of trade-offs, as measured by newly-proposed sparsity-based metrics.
Quality-Diversity, Multi-Objective Optimisation, MAP-Elites, Neuroevolution
$c$s: Theory of computation Evolutionary algorithmsccs: Applied computing Multi-criterion optimization and decision-makingccs: Computer systems organization Evolutionary robotics
@@
Figure 1. Left. Multi-Objective MAP-Elites repertoire. The feature space $\mathcal{C}\subset\mathbb{R}^{d}$ is tessellated into cells $\mathcal{C}_{i}$. A set of non-dominated solutions is stored in each cell. The aim of moqd algorithms is to fill each cell with solutions that are Pareto-optimal. Right. Overview of preference-conditioned policy gradient in the mome-p2c algorithm. By conditioning policy-gradients on updates solutions can be improved toward achieving different trade-offs of objectives (illustrated by blue arrows). By contrast, in mome-pgx, solutions are only updated to improve performance on each objective separately (illustrated by light blue arrows).
1. Introduction
@@
Over recent years, Deep Reinforcement Learning (RL) has enabled breakthroughs in mastering games (Silver et al., 2016; Mnih et al., 2013) as well as continuous control domains for locomotion (Smith et al., 2022; Cheng et al., 2024) and manipulation (OpenAI et al., 2019). These milestones have demonstrated the extraordinary potential of RL algorithms to solve specific problems. However, most approaches return only one highly-specialised solution to a single problem. In contrast, there is a growing shift in focus towards not just uncovering one single solution that achieves high rewards, but instead many solutions that exhibit different ways of doing so (Zahavy et al., 2023). Within this context, Quality-Diversity (qd) algorithms (Cully and Demiris, 2018) have emerged as one promising approach for tackling this challenge.
@@
In qd, the primary goal is to produce a variety of high-quality solutions, rather than to focus exclusively on finding the single best one. One motivation for qd algorithms is that, finding many solutions can provide availability of alternative, back-up solutions in the event that the highest-performing solution is no longer suitable. For example, in robotics, generating large collections of solutions has been shown to be helpful for addressing large simulation to reality gaps (Cully et al., 2015) and adapting to unforeseen damages (Cully et al., 2015; Allard et al., 2023). Alternatively, having multiple solutions can simply be used in order to promote innovation in the downstream task. In this context, qd has been used for creating diverse video game levels (Khalifa et al., 2018; Bhatt et al., 2022) and generating building designs (Gaier et al., 2022).
@@
Despite the growing traction of qd, most research in this field has focused on single-objective applications. However, multi-objective (mo) problems pervade many real-world domains, including engineering (Ferreira et al., 2019; Verstraeten et al., 2021), finance (Krasheninnikova et al., 2019), and drug design (Horwood and Noutahi, 2020) and many state-of-the-art mo algorithms originate from the Evolutionary Algorithm community (Zitzler and Thiele, 1998; Zitzler et al., 2001; Deb et al., 2000; Zhang and Li, 2007).
@@
Recently, Multi-Objective MAP-Elites algorithm (mome) (Pierrot et al., 2022b) marked the first attempt at bridging ideas from qd and mo optimisation. In moqd, the overarching goal is to identify a broad collection of solutions that exhibit diverse features and achieve distinct performances across multiple objectives. More specifically, given a feature space that is tessellated into cells, the aim is to find a collection of solutions within each cell which offer different trade-offs on each of the objectives (see Figure 1). As an example, consider the task of designing building sites. Within this context, it may be interesting to find different designs that vary in the number of buildings on the site. Then for each possible number of buildings, further options can be generated which present different trade-offs of ventilation and noise levels (Gaier et al., 2022). This approach equips end-users with a spectrum of viable options, thereby broadening their perspective on the array of feasible design possibilities.
@@
The mome algorithm demonstrated promising results in finding large collections of diverse solutions that balance multiple objectives. However, mome predominantly depends on random genetic variations that can cause slow convergence in large search spaces (Nilsson and Cully, 2021; Flageat et al., 2022; Pierrot et al., 2022a). This renders it less suitable for evolving neural networks with a large number of parameters.
@@
Since the inception of the mome framework, several related works exploring the domain of moqd have emerged (Janmohamed et al., 2023; Gaier et al., 2022; Dean and Cheney, 2023). Among them, mome-pgx (Janmohamed et al., 2023) builds upon the mome framework and was shown to achieve state-of-the-art performance on high-dimensional continuous control robotics tasks that can be framed as Markov Decision Processes. It uses crowding addition and selection mechanisms to encourage an even distribution of solutions on the non-dominated front and employs policy-gradient mutations for each objective function in order to drive the exploration process toward promising regions of the solution space. However, the mome-pgx approach is not without its own set of challenges. Firstly, it employs separate actor-critic networks for each objective function, which can be resource-intensive and may not scale with an increasing number of objectives. Furthermore, although using policy gradient-based updates helps with exploration in high-dimensional search spaces, the approach in mome-pgx only considers improving solutions on each objective separately. However, in the context of multi-objective problems, the goal is often not just to maximise each objective independently but rather to find solutions which offer different trade-offs among them. In this way, if end users have different preferences regarding the relative importance of each objective, they have a range of solutions to choose from.
@@
In this paper, we address the limitations of mome-pgx by introducing a new moqd algorithm: Multi-Objective Map-Elites with Preference-Conditioned Policy-Gradient and Crowding Mechanisms (mome-p2c). Rather than using a separate actor-critic framework for each objective, mome-p2c uses a single preference-conditioned actor and a single preference-conditioned critic. Similar to mome-pgx, the actor-critic framework in mome-p2c can be used to provide policy-gradient mutations which offer efficient search space exploration for high-dimensional neural-network policies. However, as illustrated in Figure 1, by conditioning the actor and critic networks on a preference, policy-gradient updates can be used to improve solutions toward achieving a given weighting over the objectives, rather than improve solutions on each objective disjointly. Moreover, using a single preference-conditioned actor-critic framework rather than one per objective also reduces training costs associated with maintaining the separate actor-critic networks of mome-pgx.
@@
We show that mome-p2c outperforms or matches the performance of mome-pgx across six robotic control moqd tasks, including newly introduced tri-objective ones (see Section 5.1). mome-p2c also outperforms mome-pgx on two newly introduced sparsity-based moqd metrics (see Section 5.3) demonstrating that it is able to attain a smoother set of trade-offs than mome-pgx. The code for mome-p2c is fully containerised and available at https://github.com/adaptive-intelligent-robotics/MOME-P2C.git.
2. Background
2.1. Quality-Diversity
@@
Quality-Diversity algorithms aim to discover collections of solutions that are both high-performing and diverse (Pugh et al., 2016). Similar to standard optimisation algorithms, a solution $\theta\in\Theta$ is assessed via a fitness function $f:\Theta\to\mathbb{R}$ that reflects its performance on the task. For example, consider the task of generating an image of a celebrity from a text prompt. In this case, the fitness of a solution could be the CLIP score (Radford et al., 2021) which measures the fidelity of an image to its caption that was used to generate it. However, an additional central component to qd algorithms, is the concept of the feature function $\Phi:\Theta\to\mathbb{R}^{d}$ that characterizes solutions in a meaningful way for the type of diversity desired (Pugh et al., 2016). The feature of a solution $\Phi(\theta_{i})$ is a vector that captures some of its notable characteristics, which is then consequently used to quantify its novelty relative to other solutions. In the image generation example, the feature could be the hair length or age of the subject in the photo (Fontaine and Nikolaidis, 2023). In this example, the qd algorithm would then aim to generate images in which the subject has a diverse range of hair lengths and ages, and which closely obey the given text prompt used to generate it.
@@
One branch of algorithms in the qd family stems from the map-elites algorithm (Mouret and Clune, 2015), which has gained prominence for its simplicity and effectiveness. map-elites operates by discretising the feature space into a grid-like structure, where each cell $C_{i}$ of the grid becomes a “niche” that can be occupied by a solution. Tessellating the feature space in this manner creates a systematic method for exploring of different niches within this space (Vassiliades et al., 2018). Each iteration of map-elites first involves selecting solutions from these niches, creating copies of them and mutating these copies to create new candidate solutions. Then, the fitness and features of the candidate solutions are evaluated, and they are added to the appropriate niches based on their fitness. If the cell corresponding to the new solution’s feature vector is unoccupied, the new solution is added to the cell. If the cell is occupied, but the evaluated solution has a higher fitness than the current occupant, it is added to the grid. Otherwise, the solution is discarded. This process continues for a fixed number of iterations, progressively populating the grid structure with an array of diverse, high-quality solutions.
@@
map-elites algorithms aim to maximise the total number of occupied cells at the end of the process and the performance of the solutions within each of them. Given a search space $\Theta$ and a feature space $\mathcal{C}$ that has been tessellated into $k$ cells $\mathcal{C}_{i}$, the map-elites objective, or QD-score (Cully and Demiris, 2018) can be formally expressed as:
(1) $\max_{\theta\in\Theta}\sum_{i=1}^{k}f(\theta_{i}),\,\,\text{where}\,\,\forall i,\Phi(\theta_{i})\in\mathcal{C}_{i}\,.$
2.2. Multi-Objective Optimisation
@@
Multi-Objective (mo) optimisation provides an approach for addressing problems that involve the simultaneous consideration of multiple, often conflicting objectives $\textbf{F}=[f_{1},\dots,f_{m}]$. In mo problems, objectives often compete with each other, meaning that improving one objective typically comes at the expense of another. For example, in engineering, improving performance might increase cost, and vice versa. To navigate this landscape, the concept of Pareto-dominance is commonly employed to establish a preference ordering among solutions. A solution $\theta_{1}$ is said to dominate another solution $\theta_{2}$ if it is equal or superior in at least one objective and not worse in any other (Hayes et al., 2022). That is, $\theta_{1}\succ\theta_{2}$, if $\forall i:f_{i}(\theta_{1})\geq f_{i}(\theta_{2})\wedge\exists j:f_{j}(\theta_{1})>f_{j}(\theta_{2})$.
@@
Solutions that are not dominated by any other solutions are termed non-dominated. Given a set of candidate solutions $\mathcal{S}$, the non-dominated solutions of this set $\theta_{i}\in S$ collectively form a non-dominated front, which represents the boundary of achievable trade-offs among objectives. The goal of mo optimisation is to find an approximation to the optimal Pareto front, i.e., the non-dominated front over the entire search space $\Theta$.
Figure 2. Two sets of solutions that form different non-dominated fronts as approximations of the true Pareto front. The outer set (marked by circles) achieves a larger hypervolume as it extends further in objective space. Likewise, its sparsity metric is higher, reflecting a more even spread of solutions. Only the outer set of solutions are considered to be non-dominated when the two sets are combined.
@@
There are two metrics, the hypervolume and sparsity metric (see Figure 2), that play pivotal roles in comprehensively assessing the quality and diversity of solutions within a Pareto front approximation (Cao et al., 2015; Hayes et al., 2022). The hypervolume $\Xi$ of a front $\mathcal{P}$, measures the volume of the objective space enclosed by a set of solutions relative to a fixed reference point $r$. This metric provides a quantitative measure of the quality and spread of solutions in the objective space and is calculated as (Hayes et al., 2022; Cao et al., 2015):
(2) $\Xi(\mathcal{P})=\lambda({\theta\in\Theta\,|\,\exists\,s\in\mathcal{P},s\succ x\succ r})$
@@
While hypervolume is widely used in the multi-objective optimisation literature as a combined measure of convergence and spread, it is also influenced by the reference point and the shape of the front. For simple linear fronts, hypervolume is maximized by evenly spaced solutions. But on inverted or irregular fronts it may rank boundary solutions more highly, and its evaluation can vary significantly with the choice of reference point, sometimes leading to counter-intuitive comparisons (Shang et al., 2021; Ishibuchi et al., 2018). We therefore acknowledge these limitations, but we follow prior work in Multi-Objective Quality-Diversity in adopting hypervolume as a standard comparative metric (Pierrot et al., 2022b; Janmohamed et al., 2023).
@@
Sparsity, by contrast, provides complementary information regarding the distribution and evenness of these solutions. It is calculated by evaluating the average nearest neighbour distance among solutions on the front, given by (Hayes et al., 2022):
(3) $S(\mathcal{P})=\frac{1}{|\mathcal{P}|-1}\sum_{j=1}^{m}\sum_{i=1}^{|\mathcal{P}|-1}(\tilde{P_{j}}(i)-\tilde{P_{j}}(i+1))^{2}$
@@
where $\tilde{P_{j}}(i)$ denotes the $i$-th solution of the list of solutions on the front $\mathcal{P}$, sorted according to the $j$-th objective and $|\mathcal{P}|$ denotes the number of solutions on the front. To ensure that the sparsity is not skewed due to different scales of each of the objectives, the objective functions must be normalised prior to calculating it. A low-sparsity metric indicates that solutions are well-dispersed through the objective space, highlighting the algorithm’s ability to provide diverse trade-off solutions. In contrast, a high-sparsity metric suggests that solutions are clustered in specific regions, potentially indicating that the algorithm struggles to explore and represent the full range of possible trade-offs.
2.3. Multi-Objective Quality-Diversity Algorithms